OpenStudy (anonymous):

If the volume of a sphere is 10293.33 cm3, what is the diameter? Use 3.14 to approximate π. Round answers to the nearest hundredth.

OpenStudy (campbell_st):

well you have \[102933.33 = \frac{4}{3} \times 3.14 \times r^3\] make r the subject... and evaluate... then double the answer for the diameter

OpenStudy (anonymous):

What would the subject be because you have all the numbers in the equation already

OpenStudy (campbell_st):

ok... start by 1 multiply both sides of the equation by 3 then divide both sides by (4 x 3.14) that will leave you with number = r^3 then take the cube root of the number to find r does that make sense..?